### 04 Apr 2026 v1

# Stochastic Stability of ACIMs for Piecewise Expanding $C^{1+\varepsilon}$ Maps

Abstract.
We prove stochastic stability of absolutely continuous invariant measures (ACIMs) for piecewise expanding $C^{1+\varepsilon}$ maps of the interval. For maps $\tau$ in the class $\mathcal{T}([0,1];s,\varepsilon)$, we consider perturbed Frobenius–Perron operators $P_{\delta}=Q_{\delta}P_{\tau}$, where $Q_{\delta}$ is a Markov smoothing operator modeling noise of intensity $\delta>0$.
In the generalized bounded variation space $BV_{1,1/p}$, we establish a Lasota–Yorke inequality uniform in $\delta$. Consequently, each $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}$, and $h_{\delta}\to h$ in $L^{1}$ as $\delta\to 0$, where $h$ is the ACIM density of $P_{\tau}$.
Our proof combines the $BV_{1,1/p}$ framework, adapted from recent ACIM existence results, with uniform quasi-compactness and perturbation theory for transfer operators. This establishes stochastic stability under minimal $C^{1+\varepsilon}$ regularity ($\varepsilon>0$), where the $C^{1}$ case is known to fail.
2020 Mathematics Subject Classification:
Primary 37A05; Secondary 37E05
1. Introduction
Piecewise expanding maps of the interval form a fundamental class of dynamical systems in one dimension. A central object in their study is an absolutely continuous invariant measure (ACIM), whose density describes the long-term statistical behavior of almost all initial conditions. A classical result of Lasota and Yorke [1] establishes the existence of ACIMs for sufficiently smooth maps with uniform expansion via what is now known as the Lasota–Yorke inequality. This approach provides control over both variation and the $L^{1}$ norm of densities and leads to strong statistical properties such as mixing and decay of correlations. Subsequent works, including those of Jakobson [2] and Bowen [3], extended these results to broader settings, typically under higher regularity assumptions.
More recently, attention has turned to systems with minimal smoothness. In particular, [4] proved the existence of ACIMs for piecewise $C^{1+\varepsilon}$ expanding maps by working in the generalized bounded variation space $BV_{1,1/p}$ defined via oscillation seminorms:
(1.1) $\|f\|_{1,1/p}=\sup_{0<r\leq 1}\frac{|\mathrm{Osc}(f,r,\cdot)|_{1}}{r^{1/p}}+\|f\|_{1}.$
This framework allows one to treat maps whose derivatives are only Hölder continuous, and the assumption of $C^{1+\varepsilon}$ regularity ($\varepsilon>0$) is essentially sharp, as $C^{1}$ regularity alone is insufficient to guarantee the existence of ACIMs.
A natural question is whether these invariant measures are stable under small perturbations. This problem, known as stochastic stability, concerns the persistence of invariant densities under the introduction of noise or small perturbations of the dynamics. It plays an important role in understanding the robustness of statistical properties of dynamical systems, particularly in applications where noise or numerical approximation is unavoidable. Classical results on stochastic stability have been established for uniformly expanding maps in spaces such as bounded variation or Hölder spaces, typically under stronger smoothness assumptions.
The purpose of this paper is to study stochastic stability of ACIMs for piecewise expanding $C^{1+\varepsilon}$ maps in the above low-regularity setting. Let $\tau$ be a map in the class $\mathcal{T}([0,1];s,\varepsilon)$ of piecewise expanding $C^{1+\varepsilon}$ maps of the interval. We consider a family of perturbed Frobenius–Perron operators of the form
(1.2) $P_{\delta}=Q_{\delta}P_{\tau},$
where $P_{\tau}$ is the Frobenius–Perron operator associated with $\tau$, and $Q_{\delta}$ is a Markov smoothing operator representing noise of intensity $\delta>0$. Working in the space $BV_{1,1/p}$, we establish a Lasota–Yorke inequality that is uniform in $\delta$. As a consequence, each operator $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}$, and we prove that $\|h_{\delta}-h\|_{L^{1}}\to 0$ as $\delta\to 0$, where $h$ is the invariant density of the unperturbed system $P_{\tau}$ in $BV_{1,1/p}$.
Our approach combines the generalized bounded variation framework with perturbation techniques for quasi-compact operators. This extends stochastic stability results to piecewise $C^{1+\varepsilon}$ maps under minimal regularity assumptions in the $BV_{1,1/p}$ setting.
The paper is organized as follows. In Section 2, we recall the definition and properties of the space $\mathrm{BV}^{1,1}_{p}$ and define the class $\mathcal{T}([0,1];s,\varepsilon)$ of piecewise expanding $C^{1+\varepsilon}$ maps. In Section 3, we define the perturbation operators and establish a uniform Lasota–Yorke inequality. Section 4 proves existence of invariant densities and stochastic stability via convergence in $L^{1}$. Section 5 discusses remarks and further directions.
2. Preliminaries
In this section, we recall the definition and basic properties of the generalized bounded variation spaces $BV_{1,1/p}$, following ([4, 5]). These spaces provide the appropriate functional setting for studying the Frobenius–Perron operator associated with piecewise expanding maps of low regularity.
Let $I=[0,1]$, and let $m$ denote the Lebesgue measure on II. For a measurable function $f:I\to\mathbb{R}$ and $r>0$, define the oscillation of $f$ at scale $r$ by
(2.1) $\mathrm{Osc}(f,r,x)=\sup\bigl\{|f(y_{1})-f(y_{2})|:y_{1},y_{2}\in(x-r,x+r)\cap I\bigr\}.$
The function $\mathrm{Osc}(f,r,\cdot)$ is measurable and describes the local variation of $f$.
For $1\leq p<\infty$, define
(2.2) $\mathrm{Osc}_{1}(f,r)=\int_{I}\mathrm{Osc}(f,r,x)\,dm(x).$
Definition 2.1.
Let $p\geq 1$. The space $BV_{1,1/p}$ consists of all functions $f\in L^{1}(I)$ such that
(2.3) $\mathrm{var}_{1,1/p}(f):=\sup_{0<r\leq 1}\frac{\mathrm{Osc}_{1}(f,r)}{r^{1/p}}<\infty.$
We equip this space with the norm
(2.4) $\|f\|_{1,1/p}=\mathrm{var}_{1,1/p}(f)+\|f\|_{1}.$
The space $(BV_{1,1/p},\|\cdot\|_{1,1/p})$ is a Banach space and is continuously embedded in $L^{1}(I)$. Moreover, the unit ball in $BV_{1,1/p}$ is relatively compact in $L^{1}(I)$.
We will use the following basic properties of oscillation.
Proposition $2.$
Let $f\in BV_{1,1/p}$ and $r>0$. Then:
  1. (1)
    $\mathrm{Osc}(f,r,\cdot)$ is lower semicontinuous and hence measurable;
  2. (2)
    $\mathrm{Osc}_{1}(f,r)$ is nondecreasing in $r$;
We now recall a key estimate that will be used repeatedly.
Proposition 2.3.
Let $\tau:I\to I$ be monotone on an interval $J\subset I$, and suppose that $|\tau^{\prime}(x)|\geq s>1$ for all $x\in J$. Then for any function $f:J\to\mathbb{R}$ and any $r>0$,
(2.5) $\mathrm{Osc}(f\circ\tau^{-1},r,y)\leq\mathrm{Osc}\left(f,\frac{r}{s},\tau^{-1}(y)\right)$
for all $y\in\tau(J)$.
This estimate reflects the contraction of inverse branches of expanding maps and plays a crucial role in establishing Lasota–Yorke type inequalities.
Finally, we define the class of maps under consideration.
Definition 2.4.
A map $\tau:I\to I$ belongs to the class $\mathcal{T}([0,1];s,\varepsilon)$ if there exists a finite partition $0=a_{0}<a_{1}<\cdots<a_{q}=1$ such that:
  1. (1)
    for each $i=1,\dots,q$, $\tau$ is monotone and $C^{1+\varepsilon}$ on $(a_{i-1},a_{i})$, extending continuously to $a_{i-1},a_{i}]$;
  2. (2)
    $|\tau^{\prime}(x)|\geq s>1$ for all $x$ where defined;
  3. (3)
    $\tau^{\prime}$ is $\varepsilon$-Hölder continuous on each interval.
Assumption 2.5 (Expansion condition (q3)).
By Theorem 3 of [4], for $\tau\in\mathcal{T}([0,1];s,\varepsilon)$, the Frobenius–Perron operator satisfies the Lasota–Yorke inequality
$\mathrm{var}_{1,1/p}(P_{\tau}f)\leq\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}$
with contraction constant $\alpha_{0}=s^{-1/p}+2s^{-1}$ and $C_{0}$ depending only on $\tau$. The hypothesis $\alpha_{0}<1/C_{1}$ of Theorem 3.5 is therefore equivalent to
$\frac{1}{s^{1/p}}+\frac{2}{s}<\frac{1}{2^{1/p}\max(1,4\|q^{\prime}\|_{\infty})},$
which holds for all sufficiently large $s$. We assume henceforth that $s$ satisfies this stronger condition.
These preliminaries will be used in the subsequent sections to analyze the perturbed operators and establish stochastic stability.
3. Perturbation Operators and Uniform Lasota–Yorke Inequality
Let $q:\mathbb{R}\to[0,\infty)$ be a $C^{1}$ function satisfying
(3.1) $\int_{\mathbb{R}}q(z)\,dz=1,\qquad\mathrm{supp}(q)\subset[-1,1].$
For $\delta\in(0,1/4)$, define the kernel $q_{\delta}(x,y)=\delta^{-1}q((x-y)/\delta)$ and the smoothing operator
$(Q_{\delta}f)(x)\;=\;\int_{I}q_{\delta}(x,y)\,f(y)\,dy,\qquad x\in I.$
Because $\mathrm{supp}(q)\subset[-1,1]$, the kernel $q_{\delta}(x,\cdot)$ is supported on $(x-\delta,\,x+\delta)$. The perturbed operator is $P_{\delta}=Q_{\delta}P_{\tau}$.
For the oscillation estimates in Lemmas 3.3 and 3.4 we use the shift-invariance of the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$. Specifically, for $x\in[\delta,\,1-\delta]$ the support of $q_{\delta}(x,\cdot)$ lies entirely within $(0,1)\subset I$, so the torus distance $d_{\mathbb{T}}$ and the Euclidean distance agree on this support, and the torus computation is identical to the interval computation. For $x$ in the boundary strips $0,\delta)\cup(1-\delta,1]$, the kernel reaches outside II; these regions contribute at most $C\|g\|_{1}$ to $\mathrm{Osc}_{1}$ and are handled by Lemma 3.2. All subsequent oscillation estimates on II follow by combining the interior torus bound with Lemma 3.2.
Proposition 3.1.
For each $\delta>0$, the operator $P_{\delta}$ is positive, preserves integrals,
$\int P_{\delta}f\,dm=\int f\,dm\quad\text{for all }f\in L^{1}(I),$
and is a contraction on $L^{1}(I)$:
$\|P_{\delta}f\|_{1}\leq\|f\|_{1}\quad\text{for all }f\in L^{1}(I).$
Moreover, if $f\geq 0$, then $\|P_{\delta}f\|_{1}=\|f\|_{1}$.
Proof.
The operators $P_{\tau}$ and $Q_{\delta}$ are positive linear operators. Moreover, $P_{\tau}$ preserves integrals, and for $Q_{\delta}$ we have
$(Q_{\delta}f)(x)=\int_{\mathbb{T}}q_{\delta}(x,y)f(y)\,dy,$
where $q_{\delta}(x,y)\geq 0$ and $\int_{\mathbb{T}}q_{\delta}(x,y)\,dx=1$ for each $y\in\mathbb{T}$ (since $\int_{\mathbb{R}}q_{\delta}(x,y)\,dx=1$ and wraparound does not occur). Thus,
$\int_{\mathbb{T}}(Q_{\delta}f)(x)\,dx=\int_{\mathbb{T}}f(y)\left[\int_{\mathbb{T}}q_{\delta}(x,y)\,dx\right]dy=\int_{\mathbb{T}}f(y)\,dy=\int_{I}f(y)\,dy.$
Thus,
$\int(Q_{\delta}f)(x)\,dx=\int f(y)\,dy,$
so $Q_{\delta}$ also preserves integrals. Hence $P_{\delta}=Q_{\delta}P_{\tau}$ is positive and preserves integrals.
To prove the $L^{1}$ contraction property, note that for any $f\in L^{1}(I)$,
$|P_{\delta}f|\leq P_{\delta}|f|,$
by positivity. Integrating and using preservation of integrals, we obtain
$\|P_{\delta}f\|_{1}=\int|P_{\delta}f|\,dm\leq\int P_{\delta}|f|\,dm=\int|f|\,dm=\|f\|_{1}.$
Finally, if $f\geq 0$, then $|P_{\delta}f|=P_{\delta}f$, so
$\|P_{\delta}f\|_{1}=\int P_{\delta}f\,dm=\int f\,dm=\|f\|_{1}.$
Lemma 3.2 (Boundary strip estimate).
For any $g\in L^{1}(I)$, $r>0$, and $0<\delta<1/4$,
$\int_{0}^{\delta}\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\,dx\;+\;\int_{1-\delta}^{1}\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\,dx\;\leq\;4\|q\|_{\infty}\|g\|_{1}.$
Proof.
For any $x\in I$, pointwise:
$\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\;\leq\;2\|Q_{\delta}g\|_{\infty}\;\leq\;\frac{2\|q\|_{\infty}}{\delta}\|g\|_{1},$
where the last bound uses $|(Q_{\delta}g)(x)|\leq\|q_{\delta}(x,\cdot)\|_{\infty}\|g\|_{1}=\delta^{-1}\|q\|_{\infty}\|g\|_{1}$. Integrating over $0,\delta)$ and $(1-\delta,1]$, each of measure $\delta$:
$\int_{0}^{\delta}\mathrm{Osc}(Q_{\delta}g,r,x)\,dx\;\leq\;\delta\cdot\frac{2\|q\|_{\infty}}{\delta}\|g\|_{1}\;=\;2\|q\|_{\infty}\|g\|_{1}.$
Summing both boundary strips gives the result. ∎
Lemma $3.$
Work on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ with $0<\delta<1/2$. Let $q_{\delta}(x,y)=\frac{1}{\delta}q\!\left(\frac{x-y}{\delta}\right)$ where $q\in C^{1}$, $\operatorname{supp}(q)\subset[-1,1]$, $q\geq 0$, $\int q=1$. Set $C:=4\|q^{\prime}\|_{\infty}$. For all $g\in L^{1}(\mathbb{T})$ and $0<r\leq\delta$,
$\mathrm{Osc}_{1}(Q_{\delta}g,\,r)\;\leq\;\frac{Cr}{\delta}\,\mathrm{Osc}_{1}(g,\,r+\delta).$
Proof.
Fix $r>0$ and $x\in\mathbb{T}$. Let $x_{1},x_{2}\in B(x,r)$, so $|x_{1}-x_{2}|\leq 2r$. The kernels $q_{\delta}(x_{i},\cdot)$ are supported on $x_{i}-\delta,\,x_{i}+\delta]\subset B(x,\,r+\delta)$ (since $|x_{i}-x|\leq r$).
Set $c:=\dfrac{1}{|B(x,r+\delta)|}\displaystyle\int_{B(x,r+\delta)}g(y)\,dy$. Since $\int q_{\delta}(x_{i},y)\,dy=1$ for each $i$,
$(Q_{\delta}g)(x_{1})-(Q_{\delta}g)(x_{2})=\int_{\mathbb{T}}\bigl(q_{\delta}(x_{1},y)-q_{\delta}(x_{2},y)\bigr)\bigl(g(y)-c\bigr)\,dy.$
The integrand is supported inside $B(x,r+\delta)$. For $y\in B(x,r+\delta)$, since both $y$ and the averaging variable $z$ lie in $B(x,r+\delta)$,
$|g(y)-c|\;=\;\left|\frac{1}{|B(x,r+\delta)|}\int_{B(x,r+\delta)}\bigl(g(y)-g(z)\bigr)\,dz\right|\;\leq\;\sup_{z\in B(x,r+\delta)}|g(y)-g(z)|\;\leq\;\mathrm{Osc}(g,\,r+\delta,\,x).$
For the kernel difference, substitute $u=(x_{1}-y)/\delta$ and apply the mean value theorem; since $\operatorname{supp}(q)\subset[-1,1]$,
$\int_{\mathbb{T}}\bigl|q_{\delta}(x_{1},y)-q_{\delta}(x_{2},y)\bigr|\,dy=\int_{\mathbb{R}}\left|q(u)-q\!\Bigl(u-\tfrac{x_{1}-x_{2}}{\delta}\Bigr)\right|du\;\leq\|q^{\prime}\|_{L^{1}}\cdot\frac{|x_{1}-x_{2}|}{\delta}\leq 2\|q^{\prime}\|_{\infty}\cdot\frac{2r}{\delta}=\frac{Cr}{\delta}$
where $\|q^{\prime}\|_{L^{1}}\leq|\operatorname{supp}(q^{\prime})|\cdot\|q^{\prime}\|_{\infty}\leq 2\|q^{\prime}\|_{\infty}$ (since $\operatorname{supp}(q)\subset[-1,1]$), and $|x_{1}-x_{2}|\leq 2r$. Combining, $\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\leq\dfrac{Cr}{\delta}\,\mathrm{Osc}(g,\,r+\delta,\,x)$. Integrating over $x\in\mathbb{T}$ gives the result. ∎
Lemma 3.4 (Translation coupling).
Work on $\mathbb{T}$. For all $g\in L^{1}(\mathbb{T})$ and $r>0$,
$\mathrm{Osc}_{1}(Q_{\delta}g,\,r)\;\leq\;\mathrm{Osc}_{1}(g,\,r+\delta).$
Proof.
Fix $x\in\mathbb{T}$ and $x_{1},x_{2}\in B(x,r)$. Set $v:=x_{2}-x_{1}$. Since $q_{\delta}(x_{2},y)=q_{\delta}(x_{1},y-v)$, the substitution u=yvu=y-v on $\mathbb{T}$ gives
$(Q_{\delta}g)(x_{2})=\int_{\mathbb{T}}q_{\delta}(x_{1},u)\,g(u+v)\,du,$
and therefore
$(Q_{\delta}g)(x_{1})-(Q_{\delta}g)(x_{2})=\int_{\mathbb{T}}q_{\delta}(x_{1},u)\bigl[g(u)-g(u+v)\bigr]\,du.$
The kernel $q_{\delta}(x_{1},\cdot)$ is supported on $x_{1}-\delta,\,x_{1}+\delta]\subset B(x,\,r+\delta)$. For each $u$ in this support: $u\in[x_{1}-\delta,\,x_{1}+\delta]\subset B(x,\,r+\delta)$, and $u+v\in[x_{2}-\delta,\,x_{2}+\delta]\subset B(x,\,r+\delta)$  (since $|(u+v)-x_{2}|=|u-x_{1}|\leq\delta$). Hence $|g(u)-g(u+v)|\leq\mathrm{Osc}(g,\,r+\delta,\,x)$. Since $\int q_{\delta}(x_{1},u)\,du=1$,
$|(Q_{\delta}g)(x_{1})-(Q_{\delta}g)(x_{2})|\;\leq\;\mathrm{Osc}(g,\,r+\delta,\,x).$
Taking the supremum over $x_{1},x_{2}\in B(x,r)$ and integrating over $x\in\mathbb{T}$ gives the result. ∎
Theorem 3.5 (Uniform Lasota–Yorke inequality).
Work on $\mathbb{T}$. Let $C_{1}:=2^{1/p}\max\bigl(1,\;4\|q^{\prime}\|_{\infty}\bigr)$. By Theorem 3 of [4], $\tau$ satisfies
$\mathrm{var}_{1,1/p}(P_{\tau}f)\;\leq\;\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}$
with $0<\alpha_{0}=s^{-1/p}+2s^{-1}<1/C_{1}$. Then there exist $\alpha\in(0,1)$ and $C>0$, independent of $\delta>0$, such that for all $f\in BV_{1,1/p}$,
$\|P_{\delta}f\|_{1,1/p}\;\leq\;\alpha\|f\|_{1,1/p}+C\|f\|_{1}.$
Proof.
Let $g:=P_{\tau}f$, so $P_{\delta}f=Q_{\delta}g$ and $\|g\|_{1}=\|f\|_{1}$. By hypothesis,
(3.2) $\mathrm{var}_{1,1/p}(g)\;\leq\;\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}.$
We bound $\mathrm{var}_{1,1/p}(Q_{\delta}g)=\sup_{r>0}r^{-1/p}\mathrm{Osc}_{1}(Q_{\delta}g,r)$ by splitting at $r=\delta$.
Large scales ($r\geq\delta$). By Lemma 3.4, $\mathrm{Osc}_{1}(Q_{\delta}g,r)\leq\mathrm{Osc}_{1}(g,r+\delta)\leq(r+\delta)^{1/p}\mathrm{var}_{1,1/p}(g)$. Since $r+\delta\leq 2r$ when $r\geq\delta$,
$r^{-1/p}\mathrm{Osc}_{1}(Q_{\delta}g,r)\;\leq\;2^{1/p}\,\mathrm{var}_{1,1/p}(g).$
Small scales ($0<r<\delta$). By Lemma 3.3,
$\mathrm{Osc}_{1}(Q_{\delta}g,r)\;\leq\;\frac{Cr}{\delta}\,\mathrm{Osc}_{1}(g,\,r+\delta)\;\leq\;\frac{Cr}{\delta}\,(r+\delta)^{1/p}\,\mathrm{var}_{1,1/p}(g).$
Since $r<\delta$, the function $r\mapsto r^{1-1/p}(r+\delta)^{1/p}\delta^{-1}$ is nondecreasing (both factors increase in $r$ for $p\geq 1$), so its supremum on $(0,\delta)$ is attained at $r\nearrow\delta$:
$\sup_{0<r<\delta}\;r^{-1/p}\mathrm{Osc}_{1}(Q_{\delta}g,r)\;\leq\;C\cdot\frac{\delta^{1-1/p}(2\delta)^{1/p}}{\delta}\,\mathrm{var}_{1,1/p}(g)\;=\;2^{1/p}C\,\mathrm{var}_{1,1/p}(g).$
Combining.
$\mathrm{var}_{1,1/p}(Q_{\delta}g)\;\leq\;C_{1}\,\mathrm{var}_{1,1/p}(g),\qquad C_{1}:=\max\!\bigl(2^{1/p},\;2^{1/p}C\bigr)=2^{1/p}\max\!\bigl(1,\;4\|q^{\prime}\|_{\infty}\bigr).$
Substituting (3.2) and using $\|P_{\delta}f\|_{1}\leq\|f\|_{1}$,
$\|P_{\delta}f\|_{1,1/p}\;\leq\;\underbrace{C_{1}\alpha_{0}}_{=:\,\alpha}\|f\|_{1,1/p}+\underbrace{(C_{1}C_{0}+1)}_{=:\,C}\|f\|_{1}.$
Since $\alpha_{0}<1/C_{1}$, we have $\alpha\in(0,1)$, and both constants are independent of $\delta$. ∎
4. Existence and Convergence of Invariant Densities
In this section, we prove the existence of invariant densities for the perturbed operators $P_{\delta}$ and establish their convergence in $L^{1}$ to the invariant density of the unperturbed system.
Theorem 4.1.
For each $\delta>0$, the operator $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}$ such that
$P_{\delta}h_{\delta}=h_{\delta},\qquad h_{\delta}\geq 0,\qquad\|h_{\delta}\|_{1}=1.$
Proof.
Let $f\in BV_{1,1/p}$ with $f\geq 0$ and $\|f\|_{1}=1$. By Theorem 3.5, there exist constants $\lambda\in(0,1)$ and $C>0$, independent of $\delta$, such that
$\|P_{\delta}g\|_{1,1/p}\leq\lambda\|g\|_{1,1/p}+C\|g\|_{1}\qquad\text{for all }g\in BV_{1,1/p}.$
Iterating and using $\|P_{\delta}^{k}f\|_{1}=\|f\|_{1}=1$ (since $f\geq 0$ and $P_{\delta}$ preserves the $L^{1}$ norm of nonnegative functions), we obtain for all $n\in\mathbb{N}$:
$\|P_{\delta}^{n}f\|_{1,1/p}\;\leq\;\lambda^{n}\|f\|_{1,1/p}+C\sum_{k=0}^{n-1}\lambda^{k}\|f\|_{1}\;\leq\;\|f\|_{1,1/p}+\frac{C}{1-\lambda}\;=:\;C^{\prime}.$
Define the Cesàro averages $h_{n}:=\frac{1}{n}\sum_{k=0}^{n-1}P_{\delta}^{k}f$. By convexity of the norm,
$\|h_{n}\|_{1,1/p}\;\leq\;\frac{1}{n}\sum_{k=0}^{n-1}\|P_{\delta}^{k}f\|_{1,1/p}\;\leq\;C^{\prime},$
so $\{h_{n}\}$ is bounded in $BV_{1,1/p}$. Since the embedding $BV_{1,1/p}\hookrightarrow L^{1}(I)$ is compact, there exist a subsequence $\{h_{n_{j}}\}$ and $h_{\delta}\in L^{1}(I)$ such that $h_{n_{j}}\to h_{\delta}$ in $L^{1}(I)$.
By lower semicontinuity of $\mathrm{var}_{1,1/p}$ with respect to $L^{1}$ convergence ([5], Proposition 2.3),
$\mathrm{var}_{1,1/p}(h_{\delta})\;\leq\;\liminf_{j\to\infty}\mathrm{var}_{1,1/p}(h_{n_{j}})\;\leq\;C^{\prime},$
so $h_{\delta}\in BV_{1,1/p}$.
To verify invariance, observe the telescoping identity
$P_{\delta}h_{n}-h_{n}\;=\;\frac{1}{n}\sum_{k=0}^{n-1}(P_{\delta}^{k+1}f-P_{\delta}^{k}f)\;=\;\frac{1}{n}(P_{\delta}^{n}f-f).$
Taking $L^{1}$ norms and using $\|P_{\delta}^{n}f\|_{1}=\|f\|_{1}=1$:
$\|P_{\delta}h_{n}-h_{n}\|_{1}\;\leq\;\frac{2}{n}\;\to\;0.$
Since $P_{\delta}:L^{1}(I)\to L^{1}(I)$ is a bounded linear operator, $h_{n_{j}}\to h_{\delta}$ in $L^{1}$ implies $P_{\delta}h_{n_{j}}\to P_{\delta}h_{\delta}$ in $L^{1}$. Combined with $\|P_{\delta}h_{n_{j}}-h_{n_{j}}\|_{1}\to 0$ and $h_{n_{j}}\to h_{\delta}$, we conclude $P_{\delta}h_{\delta}=h_{\delta}$.
Finally, since $h_{n_{j}}\to h_{\delta}$ in $L^{1}$, there is a further subsequence along which $h_{n_{j}}(x)\to h_{\delta}(x)$ almost everywhere. Since each $h_{n_{j}}\geq 0$, we get $h_{\delta}\geq 0$ a.e. Moreover,
$\|h_{\delta}\|_{1}=\lim_{j\to\infty}\|h_{n_{j}}\|_{1}=1.\qed$
Lemma 4.2.
Let $K\subset L^{1}(I)$ be relatively compact. Then
$\sup_{g\in K}\|Q_{\delta}g-g\|_{1}\to 0\quad\text{as }\delta\to 0.$
Proof.
Since $K$ is relatively compact in $L^{1}(I)$, it is totally bounded. Hence, for any $\varepsilon>0$, there exist functions $g_{1},\dots,g_{N}\in L^{1}(I)$ such that for every $g\in K$ there exists $i$ with
$\|g-g_{i}\|_{1}<\varepsilon.$
For any $g\in K$, choose such $g_{i}$. Then
$\|Q_{\delta}g-g\|_{1}\leq\|Q_{\delta}g-Q_{\delta}g_{i}\|_{1}+\|Q_{\delta}g_{i}-g_{i}\|_{1}+\|g_{i}-g\|_{1}.$
Since $Q_{\delta}$ is a Markov operator, it is a contraction on $L^{1}(I)$, so
$\|Q_{\delta}g-Q_{\delta}g_{i}\|_{1}\leq\|g-g_{i}\|_{1}<\varepsilon.$
Thus,
$\|Q_{\delta}g-g\|_{1}\leq 2\varepsilon+\|Q_{\delta}g_{i}-g_{i}\|_{1}.$
Taking the supremum over $g\in K$, we obtain
$\sup_{g\in K}\|Q_{\delta}g-g\|_{1}\leq 2\varepsilon+\max_{1\leq i\leq N}\|Q_{\delta}g_{i}-g_{i}\|_{1}.$
For each fixed $i$, $\|Q_{\delta}g_{i}-g_{i}\|_{1}\to 0$ as $\delta\to 0$ (approximation of identity). Hence,
$\max_{1\leq i\leq N}\|Q_{\delta}g_{i}-g_{i}\|_{1}\to 0.$
Therefore,
$\limsup_{\delta\to 0}\sup_{g\in K}\|Q_{\delta}g-g\|_{1}\leq 2\varepsilon.$
Since $\varepsilon>0$ is arbitrary, the result follows. ∎
Theorem 4.3 (Stochastic stability).
Assume $\tau$ admits a unique ACIM with density $h\in BV_{1,1/p}$. Let $h_{\delta}$ be an invariant density of $P_{\delta}$. Then $\|h_{\delta}-h\|_{1}\to 0$ as $\delta\to 0$.
Proof.
We define an equivalent norm on $BV_{1,1/p}$ by
$\|f\|_{\mathcal{B}}:=\mathrm{var}_{1,1/p}(f)+A\|f\|_{1},\qquad A>0\text{ chosen below},$
as the strong norm, and retain $\|\cdot\|_{L^{1}}:=\|\cdot\|_{1}$ as the weak norm.
Step 1: Uniform Lasota–Yorke inequality and operator bound. By Theorem 3.5, there exist $\alpha\in(0,1)$ and $C>0$, independent of $\delta$, such that
$\mathrm{var}_{1,1/p}(P_{\delta}f)\leq\alpha\,\mathrm{var}_{1,1/p}(f)+C\|f\|_{1},\qquad\|P_{\delta}f\|_{1}\leq\|f\|_{1}.$
Choose $A>C/(1-\alpha)$. Adding AA times the second inequality to the first:
$\|P_{\delta}f\|_{\mathcal{B}}\;\leq\;\alpha\|f\|_{\mathcal{B}}+B\|f\|_{L^{1}},\qquad B:=C+A(1-\alpha),$
uniformly in $\delta>0$. Since $\|f\|_{L^{1}}\leq A^{-1}\|f\|_{\mathcal{B}}$:
$\|P_{\delta}f\|_{\mathcal{B}}\;\leq\;M\|f\|_{\mathcal{B}},\qquad M:=\alpha+\tfrac{B}{A}=1+\tfrac{C}{A}<\infty,$
so $\sup_{\delta>0}\|P_{\delta}\|_{\mathcal{B}\to\mathcal{B}}\leq M$.
Step 2: Perturbation smallness $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}}\to 0$. Write $P_{\delta}f-P_{\tau}f=(Q_{\delta}-I)P_{\tau}f$. Since $P_{\tau}$ satisfies the same Lasota–Yorke inequality [4], it maps the unit ball of $(BV_{1,1/p},\|\cdot\|_{\mathcal{B}})$ into a set bounded in $BV_{1,1/p}$. By compactness of $BV_{1,1/p}\hookrightarrow L^{1}(I)$, the set
$K\;:=\;P_{\tau}\bigl(\{\|f\|_{\mathcal{B}}\leq 1\}\bigr)$
is relatively compact in $L^{1}(I)$. Lemma 4.2 then gives
$\sup_{\|f\|_{\mathcal{B}}\leq 1}\|(Q_{\delta}-I)P_{\tau}f\|_{1}\;=\;\sup_{g\in K}\|Q_{\delta}g-g\|_{1}\;\to\;0,$
and hence $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}}\to 0$.
Step 3: Spectral gap for $P_{\tau}$.
Quasi-compactness. By the Lasota–Yorke inequality for $P_{\tau}$ and the compact embedding $BV_{1,1/p}\hookrightarrow L^{1}(I)$, the Ionescu–Tulcea–Marinescu theorem [8] implies that $P_{\tau}$ is quasi-compact on $(BV_{1,1/p},\|\cdot\|_{\mathcal{B}})$ with $r_{\mathrm{ess}}(P_{\tau})\leq\alpha<1$.
Simplicity of eigenvalue 11. Let $g\in BV_{1,1/p}$ with $P_{\tau}g=g$. Write $g=g^{+}-g^{-}$ with $g^{\pm}\geq 0$. By positivity of $P_{\tau}$:
$P_{\tau}|g|\;\geq\;|P_{\tau}g|\;=\;|g|.$
Since $P_{\tau}$ is a Markov operator, $\int P_{\tau}|g|\,dm=\int|g|\,dm$. Since $P_{\tau}|g|\geq|g|$ almost everywhere and both functions have the same integral, it follows that
$P_{\tau}|g|\;=\;|g|\quad\text{almost everywhere.}$
Adding this to $P_{\tau}g=g$:
$2P_{\tau}g^{+}\;=\;P_{\tau}|g|+P_{\tau}g\;=\;2g^{+},$
so $P_{\tau}g^{+}=g^{+}$ a.e., and likewise $P_{\tau}g^{-}=g^{-}$ a.e. If $g^{+}\not\equiv 0$, then $g^{+}/\|g^{+}\|_{1}$ is a normalized nonnegative fixed point of $P_{\tau}$; by uniqueness of the ACIM (see [4]), $g^{+}=c^{+}h$. Similarly $g^{-}=c^{-}h$. Hence $g=(c^{+}-c^{-})h$, and $\ker(P_{\tau}-I)$ is one-dimensional.
Isolation. Since $r_{\mathrm{ess}}(P_{\tau})\leq\alpha<1$, all eigenvalues of $P_{\tau}$ on the unit circle are isolated with finite multiplicity. The eigenvalue 11 is simple by the above, so it is isolated. Its spectral projection is $\Pi f=\bigl(\int f\,dm\bigr)h$, which has rank one.
Step 4: Application of Keller–Liverani. Steps 1–3 verify the hypotheses of [6, Theorem 1]:
  1. (1)
    Uniform Lasota–Yorke: $\|P_{\delta}f\|_{\mathcal{B}}\leq\alpha\|f\|_{\mathcal{B}}+B\|f\|_{L^{1}}$, with $\alpha\in(0,1)$ and $B<\infty$ independent of $\delta$.
  2. (2)
    Perturbation smallness: $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}}\to 0$.
  3. (3)
    Spectral isolation: eigenvalue 11 of $P_{\tau}$ is isolated and simple, with rank-one spectral projection $\Pi$.
The Keller–Liverani theorem therefore yields spectral projections $\Pi_{\delta}$ (rank one for all sufficiently small $\delta$) satisfying
$\|\Pi_{\delta}-\Pi\|_{\mathcal{B}\to L^{1}}\;\to\;0\quad\text{as }\delta\to 0.$
Applying this to $h\in BV_{1,1/p}$:
$\|\Pi_{\delta}h-h\|_{1}\;\leq\;\|\Pi_{\delta}-\Pi\|_{\mathcal{B}\to L^{1}}\|h\|_{\mathcal{B}}\;\to\;0,$
and $\|\Pi_{\delta}h\|_{1}\to 1$. The unique normalized invariant density of $P_{\delta}$ is
$h_{\delta}\;=\;\frac{\Pi_{\delta}h}{\|\Pi_{\delta}h\|_{1}},$
and therefore $\|h_{\delta}-h\|_{1}\to 0$. ∎
5. Remarks and Further Directions
In this paper, we established stochastic stability of absolutely continuous invariant measures for piecewise expanding $C^{1+\varepsilon}$ maps in the framework of generalized bounded variation spaces $BV_{1,1/p}$. The key ingredient is a Lasota–Yorke inequality uniform with respect to the perturbation parameter (Thm. 3.5), which controls the spectral behavior of the perturbed Frobenius–Perron operators $P_{\delta}$.
These results demonstrate that stochastic stability persists under minimal $C^{1+\varepsilon}$ regularity assumptions ($\varepsilon>0$), extending classical results [6] that typically require higher smoothness or more restrictive function spaces such as standard BV or Hölder.
Several directions for further research remain open:
  1. (1)
    Higher dimensions: piecewise expanding maps on $\mathbb{R}^{d}$ or manifolds, adapting $BV_{1,1/p}$ via Whitney jets.
  2. (2)
    Non-uniform expansion: indifferent fixed points or critical points, combining with indifferent Lasota–Yorke estimates.
  3. (3)
    Refined spaces: $BV_{t,1/p}$ for $t>1$, exploring regularity/stability trade-offs.
These extensions could impact applications in stochastic numerics and data assimilation for low-regularity dynamics.